OpenStudy (anonymous):

What is the slope of the line passing through the points (–3, 4) and (2, –1)?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

that makes no sense to me

OpenStudy (anonymous):

A. 1 B. -1 C. 3/5 D. -3/5

OpenStudy (anonymous):

slope(m) = (y2 - y1) / (x2 - x1) (-3,4) x1 = -3 and y1 = 4 (2,-1) x2 = 2 and y2 = -1 now sub those numbers into the formula

OpenStudy (anonymous):

you still need help ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

slope = (y2 - y1) / (x2 - x1) your set of points are set up like this (x,y) make one set be set 1 and the other set be set 2.. set 1 (-3,4) ....here x1 = -3 and y1 = 4 set 2 (2,-1)...here x2 = 2 and y2 = -1 do you understand this ?

OpenStudy (anonymous):

-10/6

OpenStudy (anonymous):

slope = (y2 - y1) / (x2 - x1) x1 = -3 x2 = 2 y1 = 4 y2 = -1 we will sub now... slope = (-1-4) / (2 - (-3) slope = -5/(2 + 3) slope = -5/5 = -1 do you see how I did this ?

OpenStudy (anonymous):

yes now i under stand thank you so much

OpenStudy (anonymous):

can you help me with more?

OpenStudy (anonymous):

good to hear :) It is not really that hard as long as you label your points correctly

OpenStudy (anonymous):

I can help you with more

OpenStudy (anonymous):

What is the slope of the line passing through the points (2, 5) and (0, –4)?

OpenStudy (anonymous):

A. 9/2 B. 2/9 C.-9/2 D. -2/9

OpenStudy (anonymous):

set it up the same way... (2,5) x1 = 2 and y1 = 5 (0,-4) x2 = 0 and y2 = -4 now just sub your points into the slope formula : slope = (y2 - y1) / (x2 - x1) lets see what you can do...I will check your answer. Try to set it up.

OpenStudy (anonymous):

you going to try ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

i got 3/-4

OpenStudy (anonymous):

hold on...give me a sec to check

OpenStudy (anonymous):

slope = (-4-5) / (0 - 2) slope = -9/-2 = 9/2 show me how you are getting your answers because you are doing something wrong

OpenStudy (anonymous):

i had my numbers wrong i did (5-2) / (-4-0)

OpenStudy (anonymous):

What is the slope of the line whose equation is 4x – 6y = 12?

OpenStudy (anonymous):

you have to be sure and label your numbers correctly and put them in the right spot.

OpenStudy (anonymous):

1 more question

OpenStudy (anonymous):

if you are given an equation and you want to find the slope, put the equation in y = mx + b form, and m is your slope. Can you put the equation in y = mx + b form ?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

A. -4 B. -2/3 C. 2/3 D.6

OpenStudy (anonymous):

4x - 6y = 12 -- we want the y term by itself, so we must subtract 4x from both sides. 4x - 4x - 6y = -4x + 12 -- the 4x's cancel on the left side -6y = -4x + 12 -- now we need y by itself, so we will divide by -6 -6y/-6 = -4/-6x + 12/-6 y = 4/6x - 12/6 -- reduce y = 2/3x - 2 the slope is in the m position, so your slope is 2/3

OpenStudy (anonymous):

thank you so much i got a 100% on my test thank you thank you!!!! :)

OpenStudy (anonymous):

your very welcome...keep practicing these problems and you will ace how to do them, and soon you will be teaching others how to do them.